an unknown number of marbles. 3/8 were yellow,5/16 were green, 7/32 were blue and the remaining 6 were black, how many total marbles are there?

Answer by jorel1380(3719)

3/8n+5/16n+7/32n+6=n
12n+10n+7n+192=32n
192+29n=32n
3n=192
n=64
There were 64 marbles in all..